OpenStudy (lollylau):

Can I factor a^2 + 4?

OpenStudy (anonymous):

No, in general sense.

OpenStudy (anonymous):

With imaginary numbers: (a - 2i)(a + 2i)

OpenStudy (lollylau):

Can we avoid using imaginary numbers?

OpenStudy (slaaibak):

nope

OpenStudy (lollylau):

-.-"

OpenStudy (anonymous):

a^2 + 4 = (a + 2)^2 - 4a Does this help?

OpenStudy (lollylau):

I know that :)

OpenStudy (anonymous):

There is another formula involving trig, but I don't remember that now.

OpenStudy (lollylau):

???

OpenStudy (anonymous):

If its a^2 - 4, then, a^2 - 4 = (a + 2)(a - 2). But I assume that you've made no mistake in copying the question?

OpenStudy (lollylau):

nope

OpenStudy (anonymous):

OK, so that's wrong......

OpenStudy (anonymous):

I think you just can't factor it......

OpenStudy (lollylau):

hopeless...

OpenStudy (anonymous):

u can do it using imaginary roots

OpenStudy (anonymous):

why are you factoring it ;D

OpenStudy (anonymous):

Because that's the question?

OpenStudy (radar):

I can't factor it.........but maybe someone else can.

OpenStudy (jagatuba):

No it cannot be factored without the use of imaginary numbers.